Can someone help me with this problem?
A tree casts a shadow that measures 5 m. At the same time, a meter stick casts a shadow that is 0.4 m long. How tall is the tree?
Thanks,
Lyn

Answer by checkley71(8403)

.4/5=1/X OR .4X=5 OR X=5/.4 OR X=12.5 METERS FOR THE TREE HEIGHT